OpenStudy (anonymous):

-2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

Distribute the -2 to x and 3 in -2(x+3), and do the same for the -2 in -2(x+1)

OpenStudy (anonymous):

-2(x+3) = -2x-6 -2(x+1) = -2x-2 -2x-6 = -2x-2-4 -2x-6 = -2x-6

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

2 3 All real numbers No solution

OpenStudy (anonymous):

-2x-6=-2x-6 +6 +6 -2x = -2x divide both sides by -2 x =x all real numbers are solutions